The gym coach ordered 56 total basketballs and soccer balls for the next year. There were 10 fewer basketballs ordered than soccer balls ordered. How many of each type of balls were ordered.
Answer by unlockmath(1688)

Hello,
We can set up two equations. x be basketballs and y represents soccer balls.
x+y=56
(x+10)=y
this second equation can be rewritten as:
x-y=-10
Add this equation to the first and we get:
2x=46
x=23
there we go 23 basketball and 33 soccer balls.
Make sense?
RJ
www.math-unlock.com